OpenStudy (anonymous):

Given A(5, 5), B(3, 3), C(6, 8), and D(−3, 4), which of the following segments is longest?segment AB segment BC segment CD segment BD i got segment BCam i right!

OpenStudy (anonymous):

CD is the longest. Length of BC is sqrt(3^2+5^2)=5.83 Length of CD is sqrt(9^2+4^2)=9.84 Length of BD is sqrt(6^2+7^2)=9.22